# Interpretable Multimodal Learning for Tumor Protein-Metal Binding: Progress, Challenges, and Perspectives

### v1

@@ 3.3 Data preprocessing
Transforming raw biological data into usable features for ML models requires several modality-specific preprocessing steps. Raw data are often unstructured or semi-structured and may contain inconsistencies, noise, or missing values — factors that obscure meaningful biological patterns and degrade model performance. As such, preprocessing is essential for cleaning, normalizing, and extracting relevant signals from biological data. In tumor protein-metal binding, the scarcity of high-quality datasets requires constructing new datasets or augmenting existing ones. In these cases, reproducible and well-defined preprocessing workflows are critical to ensure the resulting data are suitable for feature extraction and downstream ML model development. To support this process, we proposed a data processing workflow for tumor protein-metal binding prediction in Fig. 2B, including the key steps and tools used to transform raw data into high-quality, structured inputs. The methods for each data modality are outlined below.
@@
Sequence data requires three main preprocessing steps to enhance data quality and prepare it for feature extraction. The first step is data filtering and cleaning. Sequences shorter than 45 to 50 residues should be removed, as they may represent incomplete segments that do not accurately reflect full proteins [46]. Then, highly similar sequences are removed to reduce the redundancy. For example, sequences can be clustered according to a certain identity threshold, usually anywhere between 25% or 90% identity, to retain only one representative from each cluster [41, 47, 48, 49]. MMseqs2 [50], HH-suite [51], CD-HIT [52], and Clustal Omega [53] can streamline the process. The next step in preprocessing is sequence alignment. This can be performed as a pairwise alignment, where a query protein sequence is compared with another to identify regions of similarity. Such alignments can help determine whether the query protein shares homology with known metal-binding proteins, suggesting a potential metal-binding function. More commonly, multiple sequence alignment (MSA) aligns a set of homologous sequences, highlighting conserved residues across the protein family — often indicative of functional or structural importance. Standard tools for performing sequence alignment include MMseqs2, BLAST [54], and T-Coffee [55]. Once aligned, sequences can be annotated with metal-binding information using known metal-binding motifs and structural information extracted from datasets. These annotated sequences form high-quality inputs for training ML models.
@@
Structure data preprocessing involves four steps. First, in data filtering and cleaning, relevant 3D structures are optimized from datasets using resolution-based filtering. For example, structures with resolutions lower than 3.0 Å are often discarded to ensure reliable geometric details [46]. Redundancy reduction is carried out by retaining only one representative structure from clusters of highly similar models, preventing the overrepresentation of a single binding conformation. Redundancy can be reduced based on sequence identity or through structural comparison using tools such as Foldseek [56]. Second, structure completion is a crucial preprocessing step that involves cleaning protein structures and imputing missing information. This process involves resolving alternate conformations, adding missing atoms (such as hydrogens) using tools like PDB2PQR [57] or Reduce, and modeling absent loops with homology modeling tools such as MODELLER [58] or structural predictions from AlphaFold. Then, the identification and annotation of metal-binding sites are performed. Metal ions and their coordinating atoms are extracted from PDB or mmCIF files by selecting HETATM records and identifying nearby residues within specific distance thresholds, typically 2.5 to 3.0 Å [46, 49]. Furthermore, geometric calculations are performed, such as determining distances, coordination numbers, and bond angles, to accurately characterize the metal coordination environment [59].
@@
Protein pocket shares several preprocessing steps with structural data, including resolution-based filtering, redundancy reduction, identifying related metal ions, and adding missing atoms. Beyond these shared steps, pocket-specific preprocessing requires defining the spatial boundaries of the protein pocket and selecting the relevant amino acid residues involved in ligand or metal coordination.
@@
Text data preprocessing in protein-metal interaction studies involves several steps to transform raw textual information into a structured and informative format suitable for computational analysis. Text preprocessing begins with extracting relevant content from datasets such as PubMed or UniProt. The raw text is cleaned to remove irrelevant characters and correct formatting inconsistencies. The next step is tokenization, where continuous text is divided into smaller units called tokens, typically words or subword segments, depending on the granularity required. Normalization techniques, such as lowercasing and stemming, are then applied to standardize the text and reduce linguistic variability. Finally, named entity recognition (NER) is employed to identify and classify key biological entities, such as protein names, metal ions, and binding sites [60, 61].
@@ 3.4 Feature extraction
Finally, we summarized the unique features that can be extracted from each data modality and presented the type of information they carry, along with their contributions to tumor protein research, as shown in Fig. 2C.
@@
Sequence-based features are extracted directly from a protein’s amino acid sequence to determine its metal-binding capabilities. First, the amino acid composition can be represented using one-hot encoding. Second, physiochemical properties, which numerically reflect each amino acid’s biochemical characteristics, offer another valuable feature. For example, each residue can be encoded by properties such as hydrophobicity, polarity, charge, molecular weight, or other experimentally derived indices [59, 46, 49]. This transforms the sequence into a vector of biochemical property profiles. In addition, evolutionary conservation profiles, which capture the frequency of amino acids observed in homologous sequences, provide additional information. Examples include position-specific scoring matrices (PSSM) [62] and hidden Markov model matrices [51], both of which are used in methods such as [63]. Finally, deep learning-based sequence embeddings have emerged as a powerful alternative to hand-crafted features. These include word2vec-like approaches such as ProtVec [64] and more advanced protein language models [65, 66]. These embeddings capture subtle sequence-level semantics and long-range dependencies often missed by traditional encoding techniques.
@@
Structure-based features are derived from the 3D structure of proteins. Secondary structure features refer to the local conformations of a protein’s backbone, which are primarily organized into alpha-helices, beta-sheets, and random coils or loops. DSSP [67] and STRIDE [68] are some of the tools used to extract these secondary structure features. These features help define the spatial orientation of metal-ligand residues, thereby identifying potential metal-binding sites. Tertiary structure features capture the complete 3D fold of a protein and include aspects such as the curvature distribution of the protein surface, contact maps that show the distances between atoms or residues in a protein chain, and the spatial convergence of residues that are distant in sequence but come together to form metal-binding pockets. Finally, quaternary structure involves organizing multiple protein subunits into a functional complex. This level of structural organization is essential for modulating protein function as it shapes interfacial interactions and cooperative binding behaviors. Features at this level include the geometry of inter-subunit interfaces, arrangements of cooperative binding sites, conformational dynamics upon ligand binding, and residue interaction networks. These various levels of structural features can be represented as 1-dimensional numerical vectors, 2-dimensional distance matrices [59, 49], or in the form of a graph using tools like Graphein [69].
@@
Protein pocket features mainly contain geometric descriptors and physicochemical features. Geometric descriptors quantitatively capture the physical characteristics of the pockets. These measurements include pocket volume, surface area, depth, and shape descriptors such as sphericity and curvature. Tools such as Fpocket [70], CASTp [71], and PocketAnchor [72] are commonly used to calculate these attributes. Physicochemical features describe the chemical environment within the pocket. For example, the distribution of hydrophobic versus hydrophilic residues, local electrostatic potential, and polarity can be analyzed by examining the amino acid composition of pocket-lining residues. These descriptors can be encoded as numerical vectors, similar to one-hot encoding and physiochemical property encodings used for sequence features. Similarly to structural data, protein pockets can be represented using advanced formats such as graph-based models. In these models, residues that form the binding pocket are treated as nodes, and their spatial proximity or physicochemical interactions are represented as edges. This approach enables the modeling of complex interaction networks that influence metal-binding behavior [49].
@@
Text features provide a valuable modality for representing descriptive biological information associated with proteins. Traditional techniques such as Bag-of-Words (BoW), Term Frequency (TF), and Inverse Document Frequency (IDF) provide a foundational approach for converting text into numerical representations. BoW captures the frequency of words without considering their order, while TF measures the relative occurrence of each word within a document. IDF adjusts these frequencies by down-weighting common terms and emphasizing rarer ones. The combination of TF and IDF, which is TF-IDF, captures the importance of terms within a document relative to the entire corpus. These simple yet effective methods help highlight relevant vocabulary and term significance in protein-related textual descriptions. Recent advances in NLP have led to the development of domain-specific language models such as SciBERT [73], BioBERT [74], and ProtST [75]. These models offer valuable initial representations or embeddings of text related to proteins, which can be used to predict their functions, including their ability to bind to metals. Additionally, relationship extraction techniques are employed to identify and structure relationships between entities. For example, a statement like “protein X binds metal Y at residue Z” can be transformed into a structured relational triplet [76]. Finally, topic modeling can identify broader themes within the text, capturing functional annotations, biological processes, or experimental conditions relevant to protein-metal interactions.
@@ 4 Multimodal learning for tumor protein-metal binding
Multimodal learning leverages the idea that complex biological entities, such as tumor proteins, can be represented through diverse data modalities, each offering a distinct perspective. In ML, these sources of information are referred to as modalities, with each modality capturing complementary aspects of the underlying biological phenomenon [77]. Integrating multiple modalities allows models to compensate for the limitations of individual data modalities and supports a more comprehensive understanding of tumor protein-metal interactions [78].
@@
In protein-metal binding prediction, four primary data modalities are particularly relevant: sequence, structure, protein pocket, and text. These modalities offer insights into different levels of protein function and behavior. For example, tumor proteins frequently harbor somatic mutations [79], which can alter their conformation and metal-binding affinity [80]. Metal binding often induces structural changes, influencing downstream biological activity [81]. Capturing both pre- and post-binding conformational states offers a more nuanced view of metal-induced functional shifts in tumor biology. The tumor microenvironment adds further complexity. Increased protein-protein interactions in this environment can promote complex formation or the emergence of new binding pockets. These dynamics highlight the need for models to integrate structural and contextual information. Textual data, in particular, should go beyond basic annotations to include a disease-specific context. For example, the silencing of MT1G in liver cancer, impairing zinc detoxification, illustrates how gene expression changes can reshape metal-binding behavior [82]. Such pathology-specific information enhances the biological relevance of predictive models and supports mechanistic interpretation.
@@
To advance multimodal learning for tumor protein-metal binding, we systematically reviewed existing models developed for tumor protein-metal binding. This section collects available open-source resources, examines commonly used combinations of data modalities, and reviews representative multimodal learning approaches, as shown in Fig. 1B.
@@ 4.1 Research advances and open resources
We reviewed existing literature to identify studies that explore the application of ML in tumor protein-metal binding. Table 3 summarizes the key characteristics of these resources based on the following criteria:
  • Programming language indicates the programming language used, reflecting its popularity and adoption within the field.
  • Modality presents the types of data modalities the model utilizes. Section 3.2 provides detailed explanations of each modality.
  • Interpretability specifies whether the model incorporates interpretability analyses beyond basic data validation or model comparison.
  • Creation and last update time reflect the recency and maintenance status, providing insight into its ongoing usability and relevance.
  • GitHub stars is a proxy of reputation and influence within the developer and research communities. This criterion is included because most relevant software is hosted and maintained on GitHub [83].
  • Paper citations reflect the academic impact and visibility of the work.
  • Tasks summarize the scope and functional capabilities of each work.
  • URL provides access to the public code repository or official website. If both are available, the code repository is prioritized.
Table 3: Summary of model characteristics for tumor protein–metal binding as of March 2025. “Localization” denotes binding site prediction. “Classification” denotes metal type classification. “Identification” denotes tumor metalloprotein identification. Models are sorted by creation date (most recent first), followed by last update time (most recent first).
Models Creation Last Programming Modality Interpretability GitHub Paper Tasks URL
update language stars citations
A. Tumor protein-metal binding models
MetalTrans
[12]
2024 2024 Python Sequence Structure Pocket Text N 1 1 Localization Classification Identification https://github.com/EduardWang/MetalTrans
MetalPrognosis
[11]
2023 2024 Python Sequence Structure Pocket Text N 5 1 Localization Classification Identification https://github.com/Jrunchang/MetalPrognosis/tree/main
MCCNN
[10]
2019 2022 Python Sequence Structure Pocket Text N N/A 65 Localization https://bitbucket.org/mkoohim/multichannel-cnn/wiki/Home
B. Protein-metal binding models
MetalNet2
[47]
2024 2025 Python Sequence N 2 1 Localization Classification Identification https://github.com/wangchulab/MetalNet2
ESMBind
[84]
2024 2024 Python Sequence Structure Pocket Y 7 1 Localization Classification Identification https://github.com/Structurebiology-BNL/ESMBind
MetalATTE
[85]
2024 2024 Python Sequence Y N/A 1 Localization Classification Identification https://huggingface.co/ChatterjeeLab/MetaLATTE
MIBPred
[86]
2023 2024 Python Sequence N 1 3 Classification Identification https://github.com/ZhangHongqi215/MIBPred
Metal3D Metal1D
[87]
2022 2024 Python Structure N 38 33 Localization Classification Identification https://github.com/lcbc-epfl/metal-site-prediction
GASS-Metal
[88]
2022 2024 C++ Python Structure Y 2 6 Localization Identification https://github.com/sandroizidoro/gassmetal-local
M-Ionic
[48]
2022 2023 Python Sequence N 4 8 Localization Classification Identification https://github.com/TeamSundar/m-ionic
MIB2
[89]
2022 2022 N/A Sequence Structure N N/A 79 Localization Classification Identification http://bioinfo.cmu.edu.tw/MIB2/
LMetalSite
[90]
2022 2022 Python Sequence Y 18 37 Localization Classification Identification https://github.com/biomed-AI/LMetalSite
MetalNet
[41]
2021 2024 Python Sequence N 18 25 Localization Classification Identification https://github.com/wangchulab/MetalNet
MEBIPRED
[91]
2021 2023 N/A Sequence N N/A 26 Localization Identification https://services.bromberglab.org/mebipred/home
MetalSiteHunter
[92]
2021 2022 Python Structure Pocket Y 4 13 Localization Classification Identification https://github.com/ClinicalAI/metal-site-hunter
BioMetAll
[93]
2020 2024 Python Structure N 11 37 Localization Classification Identification https://github.com/insilichem/biometall
MIonSite
[94]
2018 2018 Perl Sequence N 4 22 Localization Classification Identification https://github.com/LiangQiaoGu/MIonSite
Table 4: (Continued) Summary of model characteristics for tumor protein–metal binding as of March 2025. “Localization” denotes binding site prediction. “Classification” denotes metal type classification. “Identification” denotes tumor metalloprotein identification. Models are sorted by creation date (most recent first), followed by last update time (most recent first).
Models Creation Last Programming Modality Interpretability GitHub Paper Tasks URL
update language stars citations
C. Protein foundation models
Oneport
[95]
2024 2025 Python Sequence Structure Pocket Text N 13 1 Localization Classification Identification https://github.com/klemens-floege/oneprot
ProteinChat
[96]
2024 2025 Python Sequences Text N 9 3 Localization Classification Identification https://github.com/mignonjia/ProteinChat
ProTrek
[97]
2024 2024 Python Sequences Structure Text N 92 12 Localization Classification Identification https://github.com/westlake-repl/ProTrek
ProteinAligner
[98]
2024 2024 Python Sequences Structure Text N 10 0 Identification https://github.com/Alexiland/ProteinAligner
We first compiled three recent models designed for tumor protein–metal binding in Table 3A: MCCNN [10], MetalPrognosis [11], and MetalTrans [12]. MCCNN is the first deep learning model designed to predict disease-related mutations at metal-binding sites in metalloproteins, providing a novel platform for exploring the role of metal-binding disruptions in human disease. MCCNN integrates three types of data: protein sequence, 3D structural data of binding sites (including protein structure and pocket features), and metadata in the form of categorical text descriptions. Therefore, MCCNN can be considered a multimodal model incorporating sequence, structure, protein pocket, and text from four protein mutation datasets — MetalPDB [31], CancerResource2 [27], ClinVar [26], and Uniprot Humsavar [99]. MetalPrognosis [11] and MetalTrans [12] were developed to further enhance predictive performance and advance research on tumor protein–metal binding. Both models build upon the MCCNN dataset while introducing architectural improvements to boost accuracy and efficiency. MetalPrognosis eliminates the need for manual feature extraction by using sliding window sequences as input. These sequences are processed through pre-trained protein language models to extract rich semantic representations, which are then fed into a convolutional neural network to capture complex, high-level features. MetalTrans integrates multiple features using concatenation with a Transformer-based framework. MetalTrans enables comprehensive and context-aware feature extraction across modalities, ensuring exhaustive feature extraction.
To our knowledge, MCCNN, MetalPrognosis, and MetalTrans are the only ML models designed for tumor protein-metal binding prediction. Since tumor proteins are not structurally distinct from general proteins, the learning mechanisms in ML models are transferable across both domains. We broadened our analysis to include general protein-metal binding models, evaluating their potential applicability to tumor protein research. As shown in Table 3B, we have compiled 15 recent protein-metal binding prediction models that could be adapted to tumor-specific tasks through training or fine-tuning with tumor protein. For example, LMetalSite [90] and M-Ionic [48] can serve as valuable baselines for evaluating the performance of MetalPrognosis. In addition to binding prediction models, recent advances in foundation models have introduced pretraining on large-scale general datasets followed by fine-tuning for specific downstream tasks. We included four representative protein foundation models in Table 3C that potentially support tumor protein-metal binding tasks. These models use powerful protein language models to extract rich contextual representations and offer promising avenues for enhancing the diversity and depth of tumor protein–metal binding prediction.

### v2

@@ 3.3 Data preprocessing
Transforming raw biological data into usable features for ML models requires several modality-specific preprocessing steps. Raw data are often unstructured or semi-structured and may contain inconsistencies, noise, or missing values — factors that obscure meaningful biological patterns and degrade model performance. As such, preprocessing is essential for cleaning, normalizing, and extracting relevant signals from biological data. In tumor protein-metal binding, the scarcity of high-quality datasets necessitates the construction of new datasets or the augmentation of existing ones. In these cases, reproducible and well-defined preprocessing workflows are crucial to ensure that the resulting data is suitable for feature extraction and downstream ML model development. To support this process, Fig. 2B presents a typical data processing workflow for tumor protein-metal binding prediction, including the key steps and tools used to transform raw data into high-quality, structured inputs. We describe the methods for each data modality below.
@@
Sequence data requires three main preprocessing steps to enhance data quality and prepare it for feature extraction. The first step is data filtering and cleaning. Short sequences (e.g., shorter than 45 residues) are typically removed, as they may represent incomplete segments that do not accurately reflect full proteins [46]. Then, highly similar sequences are also removed to reduce redundancy. For example, sequences can be clustered according to a certain identity threshold (typically anywhere between 25% and 90%) to retain only one representative from each cluster [41, 47, 48, 49]. MMseqs2 [50], HH-suite [51], CD-HIT [52], and Clustal Omega [53] can streamline the process. The next step in preprocessing is sequence alignment. This can be performed as a pairwise alignment, where a query protein sequence is compared with another to identify regions of similarity. Such alignments can help determine whether the query protein shares homology with known metal-binding proteins, suggesting a potential metal-binding function. More commonly, multiple sequence alignment (MSA) aligns a set of homologous sequences, highlighting conserved residues across the protein family, often indicative of functional or structural importance. Standard tools for performing sequence alignment include MMseqs2, BLAST [54], and T-Coffee [55]. Once aligned, sequences can be annotated with metal-binding information using known metal-binding motifs and structural information extracted from datasets. These annotated sequences provide high-quality inputs for training ML models.
@@
Structure data preprocessing involves four steps. First, in data filtering and cleaning, relevant 3D structures are optimized from datasets using resolution-based filtering. For example, structures with resolutions lower than 3.0 Å are often discarded to ensure reliable geometric details [46]. Redundancy reduction is carried out by retaining only one representative structure from clusters of highly similar models, preventing the overrepresentation of a single binding conformation. Redundancy can also be reduced based on sequence identity or through structural comparison using tools such as Foldseek [56]. Second, structure completion is a crucial preprocessing step that involves cleaning protein structures and imputing missing information. This process involves resolving alternate conformations, adding missing atoms (such as hydrogens) using tools like PDB2PQR [57] or Reduce [58], and modeling absent loops with homology modeling tools such as MODELLER [59] or structural predictions from AlphaFold [60]. Then, the identification and annotation of metal-binding sites are performed. Metal ions and their coordinating atoms are extracted from PDB or mmCIF files by selecting HETATM records and identifying nearby residues within specific distance thresholds, typically 2.5 to 3.0 Å [46, 49]. Furthermore, geometric calculations are performed, such as determining distances, coordination numbers, and bond angles, to accurately characterize the metal coordination environment [61].
@@
Protein pocket shares the following preprocessing steps with structural data: resolution-based filtering, redundancy reduction, identification of related metal ions, and imputation of missing atoms. Beyond these shared steps, pocket-specific preprocessing requires defining the spatial boundaries of the protein pocket and selecting the relevant amino acid residues involved in ligand or metal coordination.
@@
Text data preprocessing in protein-metal interaction studies involves four steps to transform raw textual information into a structured and informative format suitable for computational analysis. Text preprocessing begins with extracting relevant content from datasets such as PubMed or UniProt. The raw text is cleaned to remove irrelevant characters and correct formatting inconsistencies. The next step is tokenization, where continuous text is divided into smaller units called tokens, typically words or subword segments, depending on the granularity required. Normalization techniques, such as lowercase and stemming, are then applied to standardize the text and reduce linguistic variability. Finally, named entity recognition (NER) is employed to identify and classify key biological entities, such as protein names, metal ions, and binding sites [62, 63].
@@ 3.4 Feature extraction
Finally, we summarize the unique features that can be extracted from each data modality and present the type of information they carry, along with their contributions to tumor protein research, as shown in Fig. 2C.
@@
Sequence-based features are extracted directly from a protein’s amino acid sequence to determine its metal-binding capabilities. First, the amino acid composition can be represented using one-hot encoding. Second, physicochemical properties, which numerically reflect each amino acid’s biochemical characteristics, offer another valuable feature. For example, each residue can be encoded by properties such as hydrophobicity, polarity, charge, molecular weight, or other experimentally derived indices [61, 46, 49]. This transforms the sequence into a vector of biochemical property profiles. Additionally, evolutionary conservation profiles capture the frequency of amino acids observed in homologous sequences, providing further information. Examples include position-specific scoring matrices (PSSM) [64] and hidden Markov model matrices [51], both of which are used in methods such as GraphBind GraphBind [65]. Finally, deep learning-based sequence embeddings have emerged as a powerful alternative to hand-crafted features. These include word2vec-like approaches such as ProtVec [66] and more advanced protein language models [67, 68]. These embeddings capture subtle sequence-level semantics and long-range dependencies often missed by traditional encoding techniques.
@@
Structure-based features are derived from the 3D structure of proteins. Secondary structure features refer to the local conformations of a protein’s backbone, which are primarily organized into alpha-helices, beta-sheets, and random coils or loops. DSSP [69] and STRIDE [70] are some of the tools used to extract these secondary structure features. These features help define the spatial orientation of metal-ligand residues, thereby identifying potential metal-binding sites. Tertiary structure features capture the complete 3D fold of a protein and include aspects such as the curvature distribution of the protein surface, contact maps that show the distances between atoms or residues in a protein chain, and the spatial convergence of residues that are distant in sequence but come together to form metal-binding pockets. Finally, quaternary structure involves organizing multiple protein subunits into a functional complex. This level of structural organization is essential for modulating protein function as it shapes interfacial interactions and cooperative binding behaviors. Features at this level include the geometry of inter-subunit interfaces, arrangements of cooperative binding sites, conformational dynamics upon ligand binding, and residue interaction networks. These various levels of structural features can be represented as one-dimensional numerical vectors, two-dimensional distance matrices [61, 49], or in the form of a graph using tools like Graphein [71].
@@
Protein pocket features mainly contain geometric descriptors and physicochemical features. Geometric descriptors quantitatively capture the physical characteristics of the pockets. These measurements include pocket volume, surface area, depth, and shape descriptors such as sphericity and curvature. Tools such as Fpocket [72], CASTp [73], and PocketAnchor [74] are commonly used to calculate these attributes. Physicochemical features describe the chemical environment within the pocket. For example, the distribution of hydrophobic versus hydrophilic residues, local electrostatic potential, and polarity can be analyzed by examining the amino acid composition of pocket-lining residues. These descriptors can be encoded as numerical vectors, similar to one-hot encoding and physicochemical property encodings used for sequence features. Similarly to structural data, protein pockets can be represented using advanced formats such as graph-based models. In these models, residues that form the binding pocket are treated as nodes, and their spatial proximity or physicochemical interactions are represented as edges. This approach enables the modeling of complex interaction networks that influence metal-binding behavior [49].
@@
Text features provide a valuable modality to represent descriptive biological information associated with proteins. Traditional techniques such as Bag-of-Words (BoW), Term Frequency (TF), and Inverse Document Frequency (IDF) provide a foundational approach for converting text into numerical representations. BoW captures the frequency of words without considering their order, while TF measures the relative occurrence of each word within a document. IDF adjusts these frequencies by down-weighting common terms and emphasizing rarer ones. The combination of TF and IDF, which is TF-IDF, captures the importance of terms within a document relative to the entire corpus. These simple yet effective methods help highlight relevant vocabulary and term significance in protein-related textual descriptions. Recent advances in NLP have led to the development of domain-specific language models such as SciBERT [75], BioBERT [76], and ProtST [77]. These models offer valuable initial representations or embeddings of text related to proteins, which can be used to predict their functions, including their ability to bind to metals. Additionally, relationship extraction techniques are employed to identify and structure relationships between entities. For example, a statement like “protein X binds metal Y at residue Z” can be transformed into a structured relational triplet [78]. Finally, topic modeling can identify broader themes within the text, capturing functional annotations, biological processes, or experimental conditions relevant to protein-metal interactions.
@@ 4 Multimodal learning for tumor protein-metal binding
Multimodal learning leverages the idea that complex biological entities, such as tumor proteins, can be represented through diverse data modalities, each offering a distinct perspective [79]. These sources of information are referred to as modalities, each capturing complementary aspects of the underlying biological phenomenon [80]. Integrating multiple modalities allows models to compensate for the limitations of individual data modalities and supports a more comprehensive understanding of tumor protein-metal interactions [81].
@@
As discussed in the previous section, four primary data modalities offer insight into different aspects of protein function and behavior. For example, tumor proteins frequently harbor somatic mutations [82], which can alter their conformation and metal-binding affinity [83]. Metal binding often induces structural changes that influence downstream biological activity [84]. Capturing both pre- and post-binding conformational states offers a more nuanced view of metal-induced functional shifts in tumor biology. The tumor microenvironment adds further complexity. Increased protein-protein interactions in this environment can promote complex formation or the emergence of new binding pockets. These dynamics highlight the need for models to integrate structural and contextual information. Textual data, in particular, should go beyond basic annotations to include a disease-specific context. For example, the silencing of MT1G in liver cancer, impairing zinc detoxification, illustrates how gene expression changes can reshape metal-binding behavior [85]. This pathology-specific information enhances the biological relevance of predictive models and supports mechanistic interpretation.
@@
To advance multimodal learning for tumor protein-metal binding, we systematically reviewed existing models developed for this purpose. This section compiles available open-source resources, examines commonly used combinations of data modalities, and reviews representative multimodal learning approaches.
Table 3: Key characteristics of machine learning models related to tumor protein–metal binding, as of May 2025. Models are sorted by creation date (most recent first) and last update time (most recent first). For the column “Modality”, “Seq.” denotes Sequence and “Struct.” denotes Structure. For the column “Tasks”, “Localization (Loc.)” denotes binding site prediction, “Classification (Class.)” denotes metal type classification, and “Identification (Ident.)” denotes tumor metalloprotein identification. The asterisk (*) in these two columns indicates that the modality or task is not explicitly stated or studied in their paper, but examining the technical content indicates that the particular modality has been used or that the particular task can be a simple extension of the studied task(s), respectively.
Model Updated Prog. Modality Interpret- GitHub Paper Tasks URL
(Created) language Seq. Struct. Pocket Text ability stars citations Loc. Class. Ident.
A. Tumor protein-metal binding models
[…]
B. Protein-metal binding models
MetalNet2 [47] 2025 (2024) Python $\checkmark$ N 3 5 $\checkmark$ $\checkmark$ https://github.com/wangchulab/MetalNet2
[…]
MetalATTE [87] 2024 (2024) Python $\checkmark$ Y N/A 1 $\checkmark$ $\checkmark$ https://huggingface.co/ChatterjeeLab/MetaLATTE
[…]
M-Ionic [48] 2023 (2022) Python $\checkmark$ Y 4 12 $\checkmark$ $\checkmark$ $\checkmark$ https://github.com/TeamSundar/m-ionic
[…]
MetalSiteHunter [94] 2022 (2021) Python $\checkmark$ $\checkmark$ Y 5 13 $\checkmark$ $\checkmark$ $\checkmark$ https://github.com/ClinicalAI/metal-site-hunter
BioMetAll [95] 2024 (2020) Python $\checkmark$ Y 12 42 $\checkmark$ $\checkmark$ $\checkmark$ https://github.com/insilichem/biometall
[…]
C. Protein foundation models
Oneport [97] 2025 (2024) Python $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ N 13 1 $\checkmark$ $\checkmark$ $\checkmark$ https://github.com/klemens-floege/oneprot
ProteinChat [98] 2025 (2024) Python $\checkmark$ $\checkmark$ Y 10 3 $\checkmark$ $\checkmark$ $\checkmark$ https://github.com/mignonjia/ProteinChat
ProTrek [99] 2025 (2024) Python $\checkmark$ $\checkmark$ $\checkmark$ N 101 13 $\checkmark$ $\checkmark$ $\checkmark$ https://github.com/westlake-repl/ProTrek
ProteinAligner [100] 2025 (2024) Python $\checkmark$ $\checkmark$ $\checkmark$ N 11 0 $\checkmark$ https://github.com/Alexiland/ProteinAligner
@@ 4.1 Research advances and open resources
We review existing literature to identify studies that explore the application of ML in tumor protein-metal binding. Table 3 summarizes the following key characteristics of these resources: Prog. language indicates the programming language used; Modality lists the data modalities used; Interpretability specifies whether the model incorporates interpretability analyses beyond basic data validation or model comparison; Creation and last update time reflects the recency and maintenance status, providing insight into its ongoing usability and relevance; GitHub stars is a proxy of reputation and influence within the developer and research communities. This criterion is included because most relevant software is hosted and maintained on GitHub [101]; Paper citations serve as a proxy of the academic impact and visibility of the work, where the citation count is based on Google Scholar; Tasks summarize the scope and functional capabilities of each work; and URL provides access to the public code repository or official website. When both are available, only the code repository is listed.
We first study three recent models designed for tumor protein–metal binding in Table 3A: multichannel convolutional neural network (MCCNN) [10], MetalPrognosis [11], and MetalTrans [12]. MCCNN is the first deep learning model designed to predict disease-related mutations at metal-binding sites in metalloproteins, providing a novel platform for exploring the role of metal-binding disruptions in human disease. MCCNN integrates three types of data: protein sequence, 3D structural data of binding sites (including protein structure and pocket features), and metadata including categorical and textual data. Therefore, MCCNN can be considered as integrating four modalities, sequence, structure, protein pocket, and text data, from four datasets of protein mutations (Tables 1 and 2): MetalPDB [30], CancerResource2 [25], ClinVar [24], and Uniprot Humsavar [102]. MetalPrognosis [11] and MetalTrans [12] are developed to further enhance predictive performance in tumor protein–metal binding. Both models were built on the MCCNN dataset, introducing architectural improvements to enhance accuracy and efficiency. MetalPrognosis eliminates the need for manual feature extraction by using sliding window sequences as input. These sequences are processed through pre-trained protein language models to extract rich semantic representations, which are then fed into a convolutional neural network to capture complex, high-level features. MetalTrans integrates multiple features via concatenation within a transformer-based framework, enabling comprehensive and context-aware feature extraction across modalities.
To the best of our knowledge, MCCNN, MetalPrognosis, and MetalTrans are the only ML models designed for tumor protein-metal binding prediction. Since tumor proteins are not structurally distinct from general proteins, the learning mechanisms in ML models are transferable across related domains. Therefore, we expand our analysis scope to include general protein-metal binding models and open their potential applicability in tumor protein research. As shown in Table 3B, we have collected 15 recent protein-metal binding prediction models that can be adapted to tumor-specific tasks through training or fine-tuning with tumor protein. For example, LMetalSite [91] and M-Ionic [48] can serve as valuable baselines for evaluating the performance of MetalPrognosis. Besides protein-metal binding prediction models, recent advances in foundation models have introduced pre-training on large-scale general datasets, followed by fine-tuning for specific downstream tasks. We include four representative protein foundation models in Table 3C that can support tumor protein-metal binding tasks. These models utilize powerful protein language models to extract rich contextual representations, offering promising avenues to enhance the diversity and depth of tumor protein–metal binding prediction.
We can draw the following insights from Table 3:
  • Research on predictive models for tumor protein-metal binding is still nascent, with the classical MCCNN model currently prevailing. This highlights the urgent need for methodological advancements in this field. In contrast, the field of protein-metal interactions has seen a proliferation of diverse models. Notably, protein foundation models based on large language models have garnered significant academic attention (e.g., high-star GitHub projects), offering new technological directions.
  • The openness of the Python ecosystem has lowered model development barriers, benefiting the growth of the research community. Current mainstream models integrate multimodal data (sequence, structure, pocket, and text), but biological authenticity modeling can be enhanced. Future models can incorporate multiprotein interaction networks and dynamic microenvironments, and leverage tumor multi-omics architectures to integrate cross-scale information, such as chemical microenvironments and radiomics.
  • Existing models primarily focus on binding site prediction. Future work can expand to predicting specific metal binding types based on site features and establishing bidirectional prediction systems for metal ligand-protein structures, which will drive the integration of metal-binding mechanism research and tumor-targeted therapy.